### v1

# The role of the irreducible mass in repetitive Penrose energy extraction processes in a Kerr black hole

Abstract
The concept of the irreducible mass ($M_{\rm irr}$) has led to the mass-energy ($M$) formula of a Kerr-Newman black hole (BH), in turn leading to an expression for its surface area $S=16\pi M_{\rm irr}^{2}$. This allowed as well the coeval identification of the reversible and irreversible transformations. The concepts of extracted and extractable energy soon followed. This new conceptual framework avoids inconsistencies recently evidenced in a repetitive Penrose process. We consider repetitive decays in the ergosphere of an initially extreme Kerr BH and show that these processes are highly irreversible. For each decay, including the BH capture of the particle falling into the horizon, the increase of the irreducible mass is much larger than the extracted energy. Correspondingly, the horizon radius increases. The energy extraction process saturates and stops when the horizon reaches the decay radius. Thus, reaching a final non-rotating Schwarzschild BH state by this accretion process is impossible. We have assessed such processes for three selected decay radii, $r=1.2M$, $1.5M$, and $1.9M$. The energy extracted until saturation is only $2.75\%$, $4.54\%$, and $2.45\%$ of the BH initial mass, while the extractable energy is reduced by $25\%$, $56\%$, and $92\%$. The BH rotational energy is mostly converted into irreducible mass, which increases to $75\%$, $83\%$, and $95\%$ of the initial BH mass. Therefore, evaluating the effect of the irreducible mass increase in the energy extraction processes in the Kerr BH ergosphere is now mandatory.
preprint: APS/123-QED
I Introduction
In a recent article [1] (hereafter Paper I), we have revisited the equations describing a single Penrose process of a test particle $\mu_{0}$ that decays into two particles $\mu_{1}$ and $\mu_{2}$ in the ergosphere of an extreme Kerr black hole (BH) with mass $M$ and spin $\hat{a}=a/M=1$. We assumed $\mu_{0}=10^{-2}M$, fulfilling the condition of being a test particle. We addressed the decay process in three radii inside the BH ergosphere: $r=1.2M$, $1.5M$, and $1.9M$. The basic assumption to describe the process was that the decay of the test particles would occur at the turning points of their trajectories. We have further shown that to identify a specific process, we could adopt two additional conditions: one that the angular momentum of the particle $\mu_{1}$ entering the BH could be assumed in all three decay processes to be equal to $p_{\phi 1}=-19.434\,\mu_{1}M$ and that the ratio of the two particles $\mu_{2}/\mu_{1}=0.78345$. This choice has the advantage, shown in Paper I, of leading to a working solution of the decay equations (see Section II therein). Still, in Paper I, we also showed that if one turns from a single Penrose process, which is linear in the BH mass and angular momentum, an attempt to apply this linear process into a sequence of repetitive events leads to a violation of energy conservation. Thus, a nonlinear treatment of the mass and angular momentum becomes necessary. The topic of this new paper is considering the nonlinear relation between the BH mass, angular momentum, and the irreducible mass [2]. The latter separates the BH transformations into reversible ones that keep constant the irreducible mass and irreversible ones in which it monotonically increases.
To approach this problem in the most general formulation, we express in Section II the general equations for the mass-energy of a Kerr Newman metric characterized by mass, charge, angular momentum, and the irreducible mass [3, 4]. We also give the upper limits to the charge and spin parameters and the expression of the horizon surface area. In addition to these classic results, we also express the inverse relation relating the irreducible mass to the BH mass, spin, and charge. In total generality, fulfilling the above limitations, we also introduce the extracted energies presented in Paper I and the novel concept of extractable energy. We recall how the maximum extractable energy cannot exceed $29.3\%$ of the BH mass of an extreme Kerr BH and $50\%$ of the mass of an extremely charged Reissner-Nordström BH. Intermediate values apply for other BH parameters of the Kerr-Newman BH.
In Section III, we extend the analytic solution of the Penrose Process, introduced in Paper I for the extreme BH, to any spin value. We show the appearance of the irreducible mass role in the first step, starting from an extreme Kerr BH, again for the same fixed assumptions (see Table 1). The fundamental result is that a Penrose process is far from being linear, and the correct theoretical approach shows it has a highly nonlinear and irreversible behavior. Even in this first example, the increase of the irreducible mass is much larger than the extracted energy.
Section IV shows a repetitive nonlinear process where we formulate the iterative process going from $n-1$ to $n$. Table 2 and Fig. 1 summarize the explicit example of decays at $r=1.2M$. We finally proceed to the conclusions in Section V.
It is clear that due to the contribution of the irreducible mass strongly increasing at each step, the energy extraction is saturated after a finite number of steps in each case. The final value of the extractable energy remains very close to the initial value, or it is reduced owing to the conversion of rotational energy into irreducible mass. In the former case, the remaining extractable energy can be extracted by a reversible electrodynamical process, as shown in the third article in this series (Rueda and Ruffini, to be submitted).
It remains to be determined how the increase of the irreducible mass will affect the emission process in any process occurring in the BH ergosphere.
II The irreducible mass
The mass-energy of the Kerr-Newman BH is [3]
$M^{2}=\left(M_{\rm irr}+\frac{e^{2}}{4M_{\rm irr}}\right)^{2}+\frac{L^{2}}{4M_{\rm irr}^{2}}$ (1)
which expresses it as a function of the BH angular momentum $L$, the charge $e$, and the irreducible mass $M_{\rm irr}$. The latter must fulfill the condition
$\frac{L^{2}+e^{2}/4}{4M_{\rm irr}^{4}}\leq 1,$ (2)
for stable configurations to exist or, equivalently,
$\hat{a}^{2}+\hat{e}^{2}\leq 1.$ (3)
where we have define the dimensionless spin $\hat{a}\equiv L/M^{2}$ and charge $\hat{e}\equiv e/M$.
The irreducible mass holds constant in reversible transformations, while it monotonically increases in irreversible transformations. Soon after the introduction of the mass-energy formula (1), Christodoulou and Ruffini [3] and Hawking [4] introduced the relation between the irreducible mass and the surface area of the BH horizon
$S=16\pi M_{\rm irr}^{2},$ (4)
which monotonically increases for irreversible transformations and is constant otherwise. For an infinitesimal change of the irreducible mass $\delta M_{\rm irr}$, the surface area increases by
$\delta S=32\pi M_{\rm irr}\delta M_{\rm irr}\ .$ (5)
Inverting Eq. (1), we can express the irreducible mass as a function of the BH mass and angular momentum
$M_{\rm irr}=M\sqrt{\frac{1-\hat{e}^{2}/2+\sqrt{1-(\hat{a}^{2}+\hat{e}^{2})}}{2}}.$ (6)
With the irreducible mass, we introduce the concept of BH extractable energy
$E_{\rm extractable}$ $\equiv M-M_{\rm irr}$
$=M\left(1-\sqrt{\frac{1+\sqrt{1-(\hat{a}^{2}+\hat{e}^{2})}-\hat{e}^{2}/2}{2}}\right),$ (7)
which is the maximum energy that can be extracted from the Kerr-Newman BH. For an extreme Kerr BH $(L=M^{2},e=0)$, the extractable energy is $E_{\rm extractable}(\hat{a}=1,\hat{e}=0)=M(1-1/\sqrt{2})\approx 0.293M$. For an extreme Reissner-Nordström BH $(L=0,e=M)$, it is $E_{\rm extractable}(\hat{a}=0,\hat{e}=1)M(1-1/2)=0.5M$.
III The first appearance of the irreducible mass increase in the Penrose process
We now specialize in the Kerr BH, i.e., $e=0$. In this case, the system of equations of the Penrose process is given by the energy and momentum conservation
$\mu_{0}\hat{E}_{0}$ $=\mu_{1}\hat{E}_{1}+\mu_{2}\hat{E}_{2},$ (8a)
$\mu_{0}\hat{p}_{\phi 0}$ $=\mu_{1}\hat{p}_{\phi 1}+\mu_{2}\hat{p}_{\phi 2},$ (8b)
$\mu_{0}\hat{p}_{r0}$ $=\mu_{1}\hat{p}_{r1}+\mu_{2}\hat{p}_{r2},$ (8c)
and the request that the decay point be a triple turning point, i.e., the vanishing of the radial momentum of the three particles therein
$\hat{p}_{r0}$ $=-\left[(\hat{E}_{0}^{2}-1)\hat{r}^{3}+2\hat{r}^{2}\right.$
$+\left.(\hat{a}^{2}\hat{E}_{0}^{2}-\hat{p}_{\phi 0}^{2}-\hat{a}^{2})\hat{r}+2(\hat{E}_{0}\hat{a}-\hat{p}_{\phi 0})^{2}\right]^{1/2},$ (9a)
$\hat{p}_{r1}$ $=-\left[(\hat{E}_{1}^{2}-1)\hat{r}^{3}+2\hat{r}^{2}\right.$
$+\left.(\hat{a}^{2}\hat{E}_{1}^{2}-\hat{p}_{\phi 1}^{2}-\hat{a}^{2})\hat{r}+2(\hat{E}_{1}\hat{a}-\hat{p}_{\phi 1})^{2}\right]^{1/2}=0,$ (9b)
$\hat{p}_{r2}$ $=\left[(\hat{E}_{2}^{2}-1)\hat{r}^{3}+2\hat{r}^{2}\right.$
$+\left.(\hat{a}^{2}\hat{E}_{2}^{2}-\hat{p}_{\phi 2}^{2}-\hat{a}^{2})\hat{r}+2(\hat{E}_{2}\hat{a}-\hat{p}_{\phi 2})^{2}\right]^{1/2}=0.$ (9c)
Assuming the incoming particle starts at rest from infinity ($\hat{E}_{0}=1$), that the decay occurs at some given $\hat{r}$, and for given $\hat{p}_{\phi 1}$ and the outgoing to infalling particle mass-ratio $\mu_{2}/\mu_{1}$, the system of equations can be solved analytically for the other variables. The solution is
$\hat{p}_{\phi 0}$ $=\frac{2\hat{a}-\sqrt{2\hat{r}\Delta}}{2-\hat{r}},$ (10a)
$\hat{E}_{1}$ $=\frac{2\,\hat{a}\,\hat{p}_{\phi 1}+\sqrt{\rho+\hat{r}^{2}\Delta\,\hat{p}_{\phi 1}^{2}}}{R},$ (10b)
$\hat{E}_{2}$ $=\frac{1}{\tilde{\mu}_{2}}-\frac{\mu_{1}}{\mu_{2}}\hat{E}_{1},$ (10c)
$\hat{p}_{\phi 2}$ $=\frac{1}{\tilde{\mu}_{2}}\hat{p}_{\phi 0}-\frac{\mu_{1}}{\mu_{2}}\hat{p}_{\phi 1},$ (10d)
$\tilde{\mu}_{1}$ $=\frac{\rho}{K+\sqrt{K^{2}-\rho^{2}[1-(\mu_{2}/\mu_{1})^{2}]}},$ (10e)
where $\tilde{\mu}_{1}=\mu_{1}/\mu_{0}$, $\tilde{\mu}_{2}=\mu_{2}/\mu_{0}$, and
$\Delta$ $=\hat{r}^{2}-2\hat{r}+\hat{a}^{2},\quad R=\hat{r}\Delta+2(\hat{r}^{2}+\hat{a}^{2}),\quad\rho=\hat{r}\,\Delta\,R$ (11a)
$K$ $=\sqrt{(\rho+\hat{r}^{2}\Delta\hat{p}_{\phi 0}^{2})(\rho+\hat{r}^{2}\Delta\hat{p}_{\phi 1}^{2})}-\hat{r}^{2}\Delta\hat{p}_{\phi 0}\hat{p}_{\phi 1}.$ (11b)
We now denote with $M_{0}$ and $L_{0}$ the initial (i.e., before the process) mass and angular momentum. Thus, its spin parameter is $\hat{a}_{0}=L_{0}/M_{0}^{2}$ and the event horizon is $r_{+,0}=M_{0}(1+\sqrt{1-\hat{a}_{0}^{2}})$. Correspondingly, the BH initial irreducible mass is $M_{\rm irr,0}=M_{0}\sqrt{\left(1+\sqrt{1-\hat{a}_{0}^{2}}\right)/2}$. After capturing the particle “1”, the BH parameters have new values
$M$ $=M_{0}+\Delta M_{0},\quad\Delta M_{0}=\hat{E}_{1}\mu_{1},$ (12a)
$L$ $=L_{0}+\Delta L_{0},\quad\Delta L_{0}=\hat{p}_{\phi 1}\mu_{1}M_{0}.$ (12b)
For the Penrose process, $\hat{E}_{1}<0$ and $\hat{p}_{\phi 1}<0$, so $\Delta M_{0}<0$ and $\Delta L_{0}<0$, hence reducing the BH mass and angular momentum, $M<M_{0}$ and $L<L_{0}$. Accordingly, we can compute the changes in the spin parameter, event horizon, and irreducible mass change by
$\Delta\hat{a}_{0}$ $=\frac{L}{M^{2}}-\frac{L_{0}}{M_{0}^{2}},$ (13a)
$\Delta r_{+,0}$ $=M(1+\sqrt{1-\hat{a}^{2}})-M_{0}(1+\sqrt{1-\hat{a}_{0}^{2}}),$ (13b)
$\Delta M_{{\rm irr},0}$ $=M\sqrt{\frac{1+\sqrt{1-\hat{a}^{2}}}{2}}-M_{0}\sqrt{\frac{1+\sqrt{1-\hat{a}_{0}^{2}}}{2}},$ (13c)
With the above, we can also define the change in the extractable energy as
$\Delta E_{{\rm extractable},0}=\Delta M_{0}-\Delta M_{{\rm irr}_{0}},$ (14)
while the actual extracted energy in the process is
$E_{{\rm extracted},0}=|\Delta M_{0}|=-\Delta M_{0}.$ (15)
For example, let us start with an extreme Kerr BH, i.e., $\hat{a}_{0}=1$. The BH horizon is $r_{+,0}=M_{0}$ ($\hat{r}_{+,0}=1$) and the irreducible mass $M_{{\rm irr},0}=M_{0}/\sqrt{2}$. The ergosphere size on the equatorial plane is $r_{{\rm ergo},0}=2M_{0}$ ($\hat{r}_{{\rm ergo},0}=2$). Furthermore, the horizon surface area is $S_{0}=16\pi M_{{\rm irr,0}}^{2}=8\pi M_{0}^{2}=25.133M_{0}^{2}$ (cf Eq. 4), and the initial extractable energy is $E_{{\rm extractable},0}=M_{0}-M_{{\rm irr,0}}=(1-\sqrt{2}/2)M_{0}=0.293M_{0}$ (cf. Eq. II).
Assuming an incoming particle that decays at $\hat{r}=1.2$, an angular momentum of the infalling particle $\hat{p}_{\phi 1}=-19.434$, and the outgoing to infalling particle mass-ratio $\mu_{2}/\mu_{1}=0.78345$, we obtain from Eqs. (10) and (11) the values of the other system parameters for a given value of the decay point. Table 1 summarizes the results for decays at $\hat{r}=1.2$, $1.5$, and $1.9$, assuming $\mu_{0}=10^{-2}M_{0}$.
For instance, in the case $\hat{r}=1.2$, the extracted energy, $E_{{\rm extracted},0}=|\Delta M_{0}|=0.00145M_{0}$, is much lower than the increase in the irreducible mass, $\Delta M_{{\rm irr},0}=0.0158M_{0}$, which has got a boost of more than ten times larger. This result demonstrates the inefficiency of the process, which converts most BH rotational energy into irreducible mass. Therefore, the process is highly irreversible. We shall return to this point below in a definite way.
Table 1: Summary of Penrose process parameters occurring inside the ergosphere of extreme Kerr BH of mass $M_{0}$ and angular momentum $L_{0}=M_{0}^{2}$. The BH captures the decay-product particle with negative energy $E_{1}=\hat{E}_{1}\mu_{1}$ and negative angular momentum $p_{\phi 1}=\hat{p}_{\phi 1}\mu_{1}M_{0}$. We list the results for the decay occurring at three selected positions: $\hat{r}=1.2$, $1.5$, and $1.9$. We choose $\hat{E}_{0}=1$, $\hat{p}_{\phi 1}=-19.434$, $\mu_{2}/\mu_{1}=0.78345$ and obtain the other parameters from the conservation equations presented in Paper I and in Section III. We adopt $\mu_{0}=10^{-2}M_{0}$ for all cases.
$\hat{r}$ $\frac{\mu_{1}}{\mu_{0}}$ $\frac{\mu_{2}}{\mu_{0}}$ $\hat{E}_{1}$ $\hat{E}_{2}$ $\hat{p}_{\phi 2}$ $\frac{M}{M_{0}}$ $\hat{a}$ $\frac{\Delta M_{{\rm irr},0}}{M_{0}}$ $\frac{E_{{\rm extracted},0}}{M_{0}}$ $\frac{E_{\rm extractable}}{M_{0}}$
$1.2$ $0.0209$ $0.0164$ $-6.936$ $69.642$ $153.872$ $0.9986$ $0.999$ $0.0158$ $0.00145$ $0.276$
$1.5$ $0.0218$ $0.0171$ $-3.520$ $62.995$ $157.485$ $0.9992$ $0.997$ $0.0245$ $0.000768$ $0.267$
$1.9$ $0.0247$ $0.0194$ $-0.501$ $52.326$ $151.737$ $0.9999$ $0.995$ $0.0328$ $0.000124$ $0.260$
IV The repetitive nonlinear process
In paper I, we studied the case of repetitive decays, according to which it would be possible to extract the entire rotational energy of an extreme Kerr BH. The process consisted of dropping, one by one, the particle of energy $E_{0}=\mu_{0}=10^{-2}M$, which decays inside the ergosphere and generates two particles with energies $E_{1}$ and $E_{2}$. The BH absorbs the particle $E_{1}$ with negative energy and angular momentum, reducing its initial mass. All decays occur at the same location, $\hat{r}$. For $\hat{r}=1.2$, It was shown that after $246$ decays, the BH reaches the final state corresponding to a non-rotating Schwarzschild BH. This result was obtained using the condition that the repetitive process is linear in the mass $M$ and the angular momentum $L$. However, Paper I showed that such a linear procedure violates energy conservation. This section shows the correct way to implement consecutive decays must consider the nonlinearity introduced by the mass-energy formula, accounting for the change of BH mass and angular momentum and, consequently, of the irreducible mass after each decay.
We start from a BH with mass $M_{0}$ and angular momentum $L_{0}$ and drop from infinity a particle with energy $E_{0}$ and angular momentum $p_{\phi 0}$ into the BH ergosphere. We fix the values $E_{0}=\mu_{0}$, $\hat{p}_{\phi 1}=-19.434$, $\mu_{2}/\mu_{1}=0.78345$, and we obtain the rest of parameters from Eqs. (10) and (11). The first decay of the sequence has been described in Section III. We can straightforwardly extend Eqs. (12) and (13) to a sequence of decays and evaluate the change of each physical quantity. The BH mass and angular momentum at the beginning of the $n$th process depends on the energy $E_{1,n-1}$ and angular momentum $p_{\phi 1}$ of the particle that the BH captured in the preceding $n-1$ process. Therefore, we have that at the $n$th process
$M_{n}$ $=M_{n-1}+\Delta M_{n-1},\quad\Delta M_{n-1}=\hat{E}_{1,n-1}\mu_{1,n-1},$ (16a)
$L_{n}$ $=L_{n-1}+\Delta L_{n-1},\quad\hat{\Delta}L_{n-1}=p_{\phi 1}\mu_{1,n-1}M_{n-1}.$ (16b)
We recall that $p_{\phi 1}$ is a given constant in our calculation strategy, as well as the energy of the incoming particle from infinity, $E_{0}=\mu_{0}=10^{-2}M_{0}$. With the above knowledge, we obtain the changes that occurred in the spin parameter, event horizon, and irreducible mass
$\Delta\hat{a}_{n-1}$ $=\frac{L_{n}}{M_{n}^{2}}-\frac{L_{n-1}}{M_{n-1}^{2}},$ (17a)
$\Delta r_{+,n-1}$ $=M_{n}\left(1+\sqrt{1-\hat{a}_{n}^{2}}\right)-M_{0}\left(1+\sqrt{1-\hat{a}_{0}^{2}}\right),$ (17b)
$\Delta M_{{\rm irr},n-1}$ $=M_{n}\sqrt{\frac{1+\sqrt{1-\hat{a}_{n}^{2}}}{2}}-M_{0}\sqrt{\frac{1+\sqrt{1-\hat{a}_{0}^{2}}}{2}}.$ (17c)
Figure 1: The sequence of $19$ non-linear repetitive Penrose energy extraction processes from an extreme Kerr BH of mass $M_{0}$ and angular momentum $L_{0}=M_{0}^{2}$ ($\hat{a}_{0}=1$). Thus, the initial horizon radius is $r_{+,0}=M_{0}$, the irreducible mass $M_{{\rm irr},0}=M_{0}/\sqrt{2}=0.707M_{0}$, and the extractable energy $E_{{\rm extractable},0}=0.293M_{0}$. In each process, the incoming particle of mass $\mu_{0}=10^{-2}M_{0}$ decays into two particles at a radius $\hat{r}=1.2$ inside the BH ergosphere. The horizontal axis indicates the decay number. Left: BH mass (red-empty circles), spin (blue-filed squares), irreducible mass (green-filled diamonds), and horizon radius (filled-orange stars). Right: extractable (purple-filled triangles) and extracted (cyan-empty diamonds) energy. The sequence stops when the horizon radius coincides with the radius where the decay happens. The final parameters are $M_{f}=0.972M_{0}$, $\hat{a}_{f}=0.980$, $r_{+,f}=1.2M_{f}$, $M_{{\rm irr},f}=0.754M_{0}$, $E_{{\rm extractable},f}=0.219M_{0}$. The extracted energy has been $E_{{\rm extracted},f}=0.0276M_{0}$. We refer to Table 2 for the numerical values of these and other quantities of the process.
Figure 1 shows the results of applying the above procedure to a sequence of processes assuming that the $n$th decay occurs at $\hat{r}_{n}=1.2M_{n}$. This example’s main point is that the self-consistent consideration of the change in the BH mass and angular momentum implies the nonlinear increase of the BH irreducible mass in each process, consistently with the BH mass-energy formula and the irreversibility implied by the BH surface area theorem.
From the above, it turns out that the process cannot be continued arbitrarily. There is a saturation point corresponding to a maximum number of decays from which the process is not allowed anymore, implying that the extraction process of the BH rotational energy ends without reaching the state of a non-rotating Schwarzschild BH.
In the numerical procedure described above, the saturation point corresponds to the decay $n_{\rm sat}$ from which complex values for the BH mass and angular momentum are obtained. This occurs when the discriminant of the square root given the value of $\hat{E}_{1}$, given by Eq. (10b), vanishes. The exact solution to that critical point is
$\hat{a}_{c}=\sqrt{\hat{r}(2-\hat{r})}.$ (18)
However, since the process is discrete, the saturation point will satisfy $\hat{a}_{f}=a_{n}(n=n_{\rm sat})\lesssim\hat{a}_{c}$. Thus, the final BH parameters at the end of the sequence of processes are $M_{f}=M_{n}(n=n_{\rm sat})$ and $L_{f}=L_{n}(n=n_{\rm sat})$. In the example of Fig. 1, where $\hat{r}_{n}=1.2$, $\hat{a}_{c}=0.9798$, saturation is reached at $n_{\rm sat}=19$. The final spin is $\hat{a}_{f}=0.9795$, confirming the above reasoning.
The physical explanation of this effect is simple. The irreducible increases its value after each decay and induces an increase in the BH surface area and the horizon radius. When the horizon radius reaches the location where the decays occur, the process stops because, on the horizon, no decay is allowed. Indeed, if one calculates where it exactly occurs by solving the equation $\hat{r}_{+}=\hat{r}$, one obtains again the relation (18). The above behavior is illustrated in Fig. 1 for decays occurring at $\hat{r}=1.2$. The complete set of values of the main parameters for each decay is given in Table 2.
The sequence of decays for the case $\hat{r}=1.5$ reaches saturation after $60$ decays, and $\hat{r}=1.9$ after $205$ decays. Figure 2 shows the final values of the BH properties reached at the saturation point, following the analogous calculation above for decay radii from $\hat{r}=1.2$ to $1.99$. The plot shows that the larger the decay radius, the lower the final spin, so the lower the rotational and extractable energy. Until now, one could think the process is extracting the rotational energy. However, the inspection of the irreducible mass shows it becomes larger and larger, approaching the BH mass and explaining why the extractable energy decreases, approaching zero. Indeed, the extracted energy lowers, also approaching zero. The above demonstrates a dramatic conclusion for the Penrose process: instead of being transferred to the ongoing particle, the BH rotational energy lost mostly converts into irreducible mass.
Figure 2: Final values of the BH properties at the saturation point, for decay radii ranging from $\hat{r}=1.2$ to $1.99$.
V Conclusions
We have considered self-consistently the change in the BH mass and angular momentum at every step in a repetitive sequence of Penrose processes in the ergosphere of an initially extreme Kerr BH. We have shown that this leads to a nonlinear increase of the BH irreducible mass and, hence, of the horizon radius. The above effect, consistent with the BH mass-energy formula, leads to the end of the repetitive process once the horizon reaches the decay radius, leading to major consequences for repetitive Penrose processes:
  1. 1.
    The energy extracted by the end of the processes is only a small fraction of the initial Kerr BH mass.
  2. 2.
    For decays near the BH horizon (e.g., $r=1.2M$), the horizon reaches the decay radius after a few tens of decays. It leads to small energy extracted ($\lesssim 3\%$), so the final BH mass, spin, and extractable energy remain close to their initial values.
  3. 3.
    The larger the decay radius, the larger the decay number until the process stops (e.g., for $r=1.9M$, after $205$ decays). The final spin and the extractable energy are largely reduced, while the energy extracted remains little, $\lesssim 3\%$ of the initial BH mass. Therefore, in these cases, the BH rotational energy has been spent mostly on increasing the irreducible mass, as shown by direct calculation of the latter (see Fig. 2).
From the above points, we conclude that the repetitive Penrose processes either change very little the BH spin and the rotational energy, leaving most of it to be extracted, or it can approach a final state of a Schwarzschild BH but by converting the BH rotational energy into irreducible mass, without real energy extraction. Therefore, it is mandatory to consider self-consistently for the effect of the irreducible mass increase in the energy extraction processes in the Kerr BH ergosphere.
As a major consequence, we are brought to the quest for alternative BH energy extraction processes that could approach reversibility as much as possible, i.e., keeping the increase of the irreducible mass as small as possible. Electrodynamical processes in the vicinity of a Kerr BH embedded in a magnetic field and surrounded by ionized matter offer a promising alternative (e.g., [5], Rueda and Ruffini, to be submitted).
Acknowledgements.
We thank Prof. D. Christodoulou and Prof. R. P. Kerr for discussions on the topic of this work.
Table 2: Main parameters for a repetitive process occurring at $r_{n}=1.2M_{n}$. The values $\hat{E}_{0}=1$, $\hat{p}_{\phi 1}=-19.344$ and $\mu_{2}/\mu_{1}=0.78345$ are kept fixed during the process. We set the incoming particle mass to $\mu_{0}=10^{-2}M_{0}$. The remaining parameters evolve in agreement with the equations governing the ballistic Penrose process.
$n$ $\frac{M_{n}}{M_{0}}$ $\frac{L_{n}}{M_{0}^{2}}$ $\frac{L_{n}}{M_{f}^{2}}$ $\frac{E_{\rm extractable},n}{M_{0}}$ $\frac{E_{\rm extracted},n}{M_{0}}$ $\frac{M_{{\rm irr},n}}{M_{0}}$ $\frac{r_{+,n}}{M_{n}}$ $\eta_{n}$ $\frac{S_{n}}{S_{0}}$ $\hat{p}_{\phi 0,n}$ $\hat{p}_{\phi 2,n}$ $\hat{\mu}_{1,n}$ $\hat{E}_{1,n}$ $\hat{\mu}_{2,n}$ $\hat{E}_{2,n}$
[…]
1 0.99855 0.99594 0.99883 0.27561 0.1450 0.72294 1.0483 14.4958 1.0453 2.1127 153.83 2.0901 6.9356 1.6374 69.923
2 0.99710 0.99189 0.99767 0.26839 0.2899 0.72871 1.0682 14.4960 1.0620 2.1213 154.90 2.0813 6.9651 1.6306 70.219
3 0.99565 0.98787 0.99651 0.26284 0.4349 0.73281 1.0834 14.4962 1.0740 2.1301 156.01 2.0723 6.9954 1.6235 70.523
4 0.99420 0.98386 0.99537 0.25817 0.5799 0.73603 1.0962 14.4965 1.0835 2.1393 157.16 2.0632 7.0266 1.6164 70.835
5 0.99275 0.97987 0.99423 0.25407 0.7248 0.73869 1.1073 14.4967 1.0913 2.1488 158.35 2.0539 7.0586 1.6091 71.156
6 0.99130 0.97589 0.99309 0.25036 0.8698 0.74094 1.1173 14.4969 1.0980 2.1587 159.59 2.0444 7.0917 1.6017 71.487
7 0.98985 0.97194 0.99197 0.24697 1.0148 0.74288 1.1265 14.4972 1.1037 2.1691 160.88 2.0347 7.1258 1.5940 71.829
8 0.98840 0.96800 0.99085 0.24383 1.1598 0.74457 1.1349 14.4974 1.1088 2.1799 162.23 2.0247 7.1611 1.5862 72.183
9 0.98695 0.96409 0.98975 0.24090 1.3048 0.74606 1.1428 14.4976 1.1132 2.1913 163.65 2.0144 7.1978 1.5782 72.551
10 0.98550 0.96019 0.98865 0.23814 1.4498 0.74736 1.1502 14.4979 1.1171 2.2033 165.15 2.0039 7.2361 1.5699 72.934
11 0.98405 0.95632 0.98757 0.23553 1.5948 0.74852 1.1572 14.4981 1.1206 2.2159 166.73 1.9929 7.2761 1.5613 73.335
12 0.98260 0.95247 0.98650 0.23306 1.7398 0.74954 1.1638 14.4984 1.1236 2.2294 168.42 1.9815 7.3182 1.5524 73.757
13 0.98115 0.94864 0.98544 0.23071 1.8848 0.75044 1.1700 14.4987 1.1263 2.2439 170.23 1.9696 7.3628 1.5431 74.203
14 0.97970 0.94484 0.98440 0.22847 2.0299 0.75123 1.1760 14.4989 1.1287 2.2597 172.18 1.9570 7.4104 1.5332 74.681
15 0.97825 0.94106 0.98337 0.22633 2.1749 0.75192 1.1816 14.4992 1.1308 2.2769 174.33 1.9436 7.4619 1.5227 75.196
16 0.97680 0.93731 0.98236 0.22429 2.3199 0.75251 1.1870 14.4995 1.1325 2.2962 176.74 1.9291 7.5185 1.5113 75.764
17 0.97535 0.93360 0.98138 0.22235 2.4650 0.75300 1.1921 14.4998 1.1340 2.3184 179.51 1.9129 7.5825 1.4986 76.405
18 0.97390 0.92991 0.98043 0.22050 2.6100 0.75340 1.1969 14.5001 1.1352 2.3455 182.88 1.8940 7.6586 1.4838 77.168
19 0.97245 0.92628 0.97951 0.21876 2.7551 0.75369 1.2014 14.5004 1.1361 2.3831 187.54 1.8691 7.7611 1.4644 78.195